### v3

# RM-R1: Reward Modeling as Reasoning

Abstract
Reward modeling is essential for aligning large language models with human preferences through reinforcement learning from human feedback. To provide accurate reward signals, a reward model (RM) should stimulate deep thinking and conduct interpretable reasoning before assigning a score or a judgment. Inspired by recent advances of long chain-of-thought on reasoning-intensive tasks, we hypothesize and validate that integrating reasoning capabilities into reward modeling significantly enhances RM’s interpretability and performance. To this end, we introduce a new class of generative reward models – Reasoning Reward Models (ReasRMs) – which formulate reward modeling as a reasoning task. We propose a reasoning-oriented training pipeline and train a family of ReasRMs, RM-R1. RM-R1 features a chain-of-rubrics (CoR) mechanism – self-generating sample-level chat rubrics or math/code solutions, and evaluating candidate responses against them. The training of RM-R1 consists of two key stages: (1) distillation of high-quality reasoning chains and (2) reinforcement learning with verifiable rewards. Empirically, our models achieve state-of-the-art performance across three reward model benchmarks on average, outperforming much larger open-weight models (e.g., INF-ORM-Llama3.1-70B) and proprietary ones (e.g., GPT-4o) by up to 4.9%. Beyond final performance, we perform thorough empirical analyses to understand the key ingredients of successful ReasRM training. To facilitate future research, we release six ReasRM models along with code and data at https://github.com/RM-R1-UIUC/RM-R1.
1 Introduction
Reward models (RMs) play a critical role in large language model (LLM) post-training, particularly in reinforcement learning with human feedback (RLHF) [4, 24], where they serve as scalable proxies for human evaluators. Existing research on reward modeling can be broadly classified into two categories: (1) scalar-based reward models (ScalarRM)  [20] and (2) generative reward models (GenRM)  [40]. Scalar-based approaches frame reward modeling as a classification problem, typically training a sequence classifier on top of a language model. In contrast, generative approaches retain the original language model decoding head and leverage the model’s generative abilities to produce free-form pairwise judgments. While scalar-based methods are direct and often effective, they are opaque, offering no intermediate reasoning steps to justify the model’s decisions. This lack of transparency may limit their capacity to handle more challenging, reasoning-intensive preference tasks. On the other hand, although generative methods provide greater transparency, their reasoning is often superficial and unhelpful for reliable judgment, leading to suboptimal performance [7, 22].
Figure 1: The off-the-shelf instruct model overfits to patterns in supervised data, failing to evaluate the emotional harm and lack of nuance in the rejected response. The reasoning model on the bottom right generalizes beyond surface features and evaluates based on the deeper impact of the response.
In real-world decision-making scenarios, accurate and grounded reward modeling often requires jointly conducting reasoning and reward assignment. This is because preference judgments inherently involve multifaceted cognitive considerations, such as inferring a judge’s latent evaluation criteria [5], navigating trade-offs among multiple criteria [23], and simulating potential consequences [33], all of which necessitate extensive reasoning. Our example in Figure 1 illustrates such an example, where a correct preference judgement requires accurate perception of the question, understanding of the corresponding evaluation rubrics with convincing arguments – closely mirroring how humans approach grading tasks. Motivated by these observations, we explore the following central question:
Can we cast reward modeling as a reasoning task?
In this work, we unleash the reasoning potential of RMs and propose a new class of models: Reasoning Reward Models (ReasRMs). Different from standard GenRMs, ReasRMs emphasize leveraging long and coherent reasoning chains during the judging process to enhance the model’s ability to assess and distinguish complex outputs accurately. We validate that integrating long reasoning chains during the judging process significantly enhances downstream reward model performance. We explore several strategies for adapting instruction-tuned language models into logically coherent ReasRMs. Notably, we find that solely applying reinforcement learning with verifiable rewards (RLVR) [12] in reward modeling does not fully realize the model’s reasoning capabilities. We also observe that plain chain-of-thought (CoT) reasoning falls short at perceiving the fine-grained distinction across different question types.
Through a series of studies, we design a training pipeline that introduces reasoning distillation prior to RLVR, ultimately resulting in the development of RM-R1. To fully elicit the reasoning capability of RM-R1 for reward modeling, we design a Chain-of-Rubrics (CoR) process. Specifically, the model categorizes the input sample into one of two categories: chat or reasoning. For chat tasks, the model generates a set of evaluation rubrics, justifications for the rubrics, and evaluations tailored to the specific question. For reasoning tasks, correctness is the most important and generally preferred rubrics, so we directly let the model first solve the problem itself before evaluating and picking the preferred response. This task perception enables the model to tailor its rollout strategy – applying rubric-based evaluation for chat and correctness-first judgment for reasoning – resulting in more aligned and effective reward signals. In addition, we explore how to directly adapt existing reasoning models into reward models. Since these models have already undergone substantial reasoning-focused distillation, we fine-tune them using RLVR without additional distillation stages. Based on our training recipes, we produce RM-R1 models ranging from 7B to 32B.
Empirically, RM-R1 models consistently yield highly interpretable and coherent reasoning traces. On average, RM-R1 achieves state-of-the-art performance on RewardBench [17], RM-Bench [21], and RMB [43], outperforming 70B, 340B, GPT-4o, and Claude models by up to 4.9%. Beyond final performance, we conduct extensive empirical analyses of RM-R1, including ablations of our training recipes, studies of its scaling effects, comparisons with non-reasoning baselines, detailed case studies, and training dynamics. In summary, our main contributions are as follows:
  • We demonstrate that reasoning abilities are crucial for reward models, and propose to formulate reward modeling as a reasoning process to enhance interpretability and accuracy.
  • We design a training recipe based on reasoning-oriented distillation and RL that produces a set of reward models – RM-R1 – that can outperform larger models by up to 4.9% on average.
  • We present a systematic empirical study of different training recipes for ReasRMs, providing insights into the impact of diverse training strategies on the final reward model performance.
2 RM-R1
Figure 2: Training pipeline of RM-R1. Starting from an instruct model (GenRM), RM-R1 training involves two stages: Distillation and Reinforcement Learning (RL). In the Distillation stage, we use high-quality synthesized data to bootstrap RM-R1’s reasoning ability. In the RL stage, RM-R1’s reasoning ability for reward modeling is further strengthened. After distillation, a GenRM evolves into a ReasRM. RM-R1 further differentiates itself by being RL finetuned on preference data.
Figure 2 presents the overall training pipeline of RM-R1, which consists of two stages: reasoning distillation and reinforcement learning. (1) Reasoning Distillation: Starting from an off-the-shelf instruction-tuned model (e.g., Qwen-2.5-14B-Instruct), we further train the model using synthesized high-quality reasoning traces. This stage equips RM-R1 with essential reasoning capabilities required for effective reward modeling. (2) Reinforcement learning: While distillation is effective for injecting reasoning patterns, distilled models often overfit to specific patterns in the training data, limiting their generalization ability [9]. To overcome this limitation, we introduce a reinforcement learning phase that further optimizes the model, resulting in the final version of RM-R1.
2.1 Task Definition
Given a preference dataset:
$\mathcal{D}=\{(x^{(i)},y_{a}^{(i)},y_{b}^{(i)},l^{(i)})\}_{i=1}^{N},$ (1)
where $x$ is a prompt, $y_{a}$ and $y_{b}$ are two different responses for $x$, and $l\in\{a,b\}$ is the ground truth label that indicates the preferred response. We define the generative reward modeling task as follows:
Let $r_{\theta}$ denote a generative reward model parameterized by $\theta$. For each data sample, $r_{\theta}$ generates a textual judgment $j$ consisting of ordered tokens $j=(j_{1},j_{2},\dots,j_{T})$, modeled by:
$r_{\theta}(j|x,y_{a},y_{b})=\prod_{t=1}^{T}r_{\theta}(j_{t}|x,y_{a},y_{b},j_{<t}).$ (2)
Note that $j$ contains $r_{\theta}$’s prediction of the preferred response $\hat{l}\subset j$. The overall objective is:
$\max_{r_{\theta}}\mathbb{E}_{(x,y_{a},y_{b},l)\sim\mathcal{D},\hat{l}\sim r_{\theta}(j\mid x,y_{a},y_{b})}\left[\mathbbm{1}(\hat{l}=l)\right].$ (3)
2.2 Reasoning Distillation for Reward Modeling
For an instruction-tuned model (e.g., Qwen-2.5-14b-instruct [37]), it is quite intuitive to turn it into a GenRM simply by prompting. However, without fine-tuning on reward modeling reasoning traces, these models may struggle to conduct consistent judgments. To bootstrap its reasoning potential, we start with training an instruction-tuned model with long reasoning traces synthesized for reward modeling. Specifically, we sample $M$ data samples from $\mathcal{D}$ and denote it as $\mathcal{D}_{\rm sub}$. Given a data sample $(x^{(i)},y_{a}^{(i)},y_{b}^{(i)},l^{(i)})\in\mathcal{D}_{\rm sub}$, we ask an “oracle” model like o3 or claude-3-7-sonnet to generate its structured reasoning trace $r^{(i)}$ justifying why $y_{l}^{(i)}$ is chosen as the preferred response of $x^{(i)}$. We then construct the reasoning trace ground truth:
$y_{\mathrm{trace}}^{(i)}=r^{(i)}\oplus l^{(i)},$ (4)
where $\oplus$ denotes string concatenation. Given all the synthesized reasoning traces $r^{(i)}$, the final distillation dataset is defined as:
$\mathcal{D}_{\mathrm{distill}}=\{(x^{(i)},y_{\mathrm{trace}}^{(i)})\}_{i=1}^{M}.$ (5)
Formally, the objective of distillation is to adjust $\theta$ to maximize the likelihood of generating the desired reasoning trace and picking the response $y$ given the prompt $x$. We minimize the negative log-likelihood (NLL) loss:
$\mathcal{L}_{\mathrm{distill}}(\theta)=-\sum_{(x,y)\in\mathcal{D}_{\mathrm{distill}}}\sum_{t\in[|y|]}\log r_{\theta}\left(y_{t}\mid x,y_{<t}\right),$ (6)
where $y_{<t}=(y_{1},y_{2},...,y_{t-1})$ denotes the sequence of tokens preceding position $t$. More details of generating high-quality reasoning chains are included in Appendix B.
2.3 RL Training
Although distillation is a proper way to turn a general generative model into a GenRM, it often suffers from overfitting to certain patterns and constrains the model’s ability to generalize its reasoning abilities for critical thinking [9, 31] , which is essential for reward modeling. To address this, we propose to integrate RL as a more powerful learning paradigm to enhance the model’s ability to conduct reasoning-based rewarding. Training a policy model using RL has been widely seen in the preference optimization phase of LLM post-training [24], and it is quite natural to extend this paradigm for training a ReasRM. To be specific, we directly treat our reward model $r_{\theta}(j\mid x,y_{a},y_{b})$ as if it is a policy model:
$\max_{r_{\theta}}\mathbb{E}_{(x,y_{a},y_{b},l)\sim\mathcal{D},\hat{l}\sim r_{\theta}(j\mid x,y_{a},y_{b})}\left[\mathcal{R}(x,j)\right]-\beta\mathbb{D}_{\mathrm{KL}}\left(r_{\theta}\|r_{\mathrm{ref}}\right),$ (7)
where $r_{\mathrm{ref}}$ is the reference reward model. In practice, we use the checkpoint before RL training as $r_{\mathrm{ref}}$, and that means $r_{\mathrm{ref}}$ could be an off-the-shelf LLM or the LLM obtained after the distillation step in Section 2.2. $\mathcal{R}(x,j)$ is the reward function, and $\mathbb{D}_{\mathrm{KL}}$ is KL-divergence. The $x$ denotes input prompts drawn from the preference data $\mathcal{D}$. The $j$ indicates the text generated by the reward model, which includes the reasoning trace and final judgement $\hat{l}$. In practice, we use Group Relative Policy Optimization (GRPO) [28] to optimize the objective in Equation 7, the details of which can be find in Appendix C.
2.3.1 Chain-of-Rubrics (CoR) Rollout
Chain-of-Rubrics (CoR) Rollout for Instruct Models Please act as an impartial judge and evaluate the quality of the responses provided by two AI Chatbots to the Client’s question displayed below. First, classify the task into one of two categories: <type> Reasoning </type> or <type> Chat </type>. - Use <type> Reasoning </type> for tasks that involve math, coding, or require domain knowledge, multi-step inference, logical deduction, or combining information to reach a conclusion. - Use <type> Chat </type> for tasks that involve open-ended or factual conversation, stylistic rewrites, safety questions, or general helpfulness requests without deep reasoning. If the task is Reasoning:
1. Solve the Client’s question yourself and present your final answer within <solution> … </solution> tags.
2. Evaluate the two Chatbot responses based on correctness, completeness, and reasoning quality, referencing your own solution. 3. Include your evaluation inside <eval> … </eval> tags, quoting or summarizing the Chatbots using the following tags: - <quote_A> … </quote_A> for direct quotes from Chatbot A - <summary_A> … </summary_A> for paraphrases of Chatbot A - <quote_B> … </quote_B> for direct quotes from Chatbot B - <summary_B> … </summary_B> for paraphrases of Chatbot B 4. End with your final judgment in the format: <answer>[[A]]</answer> or <answer>[[B]]</answer> If the task is Chat:
1. Generate evaluation criteria (rubric) tailored to the Client’s question and context, enclosed in <rubric>…</rubric> tags.
2. Assign weights to each rubric item based on their relative importance. 3. Inside <rubric>, include a <justify>…</justify> section explaining why you chose those rubric criteria and weights. 4. Compare both Chatbot responses according to the rubric. 5. Provide your evaluation inside <eval>…</eval> tags, using <quote_A>, <summary_A>, <quote_B>, and <summary_B> as described above. 6. End with your final judgment in the format: <answer>[[A]]</answer> or <answer>[[B]]</answer>
Figure 3: The system prompt used for RM-R1 rollout.
To facilitate the distilled models to proactively generate effective reasoning traces, we design a system prompt as shown in Figure 3 during rollout. Intuitively, reward modeling for general domain (e.g., chat, safety, etc.) and reasoning domain (e.g., math, code, etc.) should focus on different angles. For example, for the chat domain, we may care more about some aspects that can be expressed in textual rubrics (e.g., be polite), yet for the reasoning domain, we usually care more about logical coherence and answer correctness. Based on this intuition, we instruct ${r_{\theta}}$ to classify each preference data sample $\{(x,y_{c},y_{r})\}$ into one of the two <type>: Chat or Reasoning. For each <type>, we prompt ${r_{\theta}}$ to carry out the behavior corresponding to that type step by step: For reasoning tasks, we ask ${r_{\theta}}$ to solve $x$ on its own. During the <eval> phase, $r_{\theta}$ compares $y_{c}$ and $y_{r}$ conditioned on its own </solution> and selects an <answer>. Regarding the Chat type, we instead ask ${r_{\theta}}$ to think about and justify the <rubric> for grading the chat quality (including safety).
2.3.2 Reward Design
Rule-based reward mechanisms have demonstrated strong empirical performance to facilitate reasoning [12]. In our training, we further simplify the reward formulation and merely focus on the correctness-based component, in line with prior works [28, 18].
Formally, our reward is defined as follows:
$\mathcal{R}(x,j|y_{a},y_{b})=\begin{cases}1&\text{if\ }\hat{l}=l,\\ -1&\text{otherwise.}\end{cases}$ (8)
where $\hat{l}$ is extracted from $j$, wrapped between the <answer> and </answer> tokens. We have also tried adding the format reward to the overall reward, but found that the task performance does not have a significant difference. The rationale behind only focusing on correctness is that the distilled models have already learned to follow instructions and format their responses properly.
3 Experiments
3.1 Experimental Setup
We evaluate RM-R1 on three primary benchmarks: RewardBench [17], RM-Bench [21], and RMB [43]. Our training set utilizes a cleaned subset of Skywork Reward Preference 80K [20], 8K examples from Code-Preference-Pairs, and the full Math-DPO-10K [16] dataset. For baselines, we compare RM-R1 with RMs from three main categories: ScalarRMs, GenRMs, and ReasRMs. Further details on the benchmarks, dataset construction, and specific baseline models are provided in Appendix D.
Table 1: The performance comparison between best-performing baselines. Bold numbers indicate the best performance, Underlined numbers indicate the second best. The DeepSeek-GRM models are not open-weighted, so we use the numbers on their tech report. The more detailed numbers on RewardBench, RM-Bench, and RMB are in Appendix Table 6,  Table 7, and Table 8
Models RewardBench RM-Bench RMB Average
ScalarRMs
SteerLM-RM-70B 88.8 52.5 58.2 66.5
Eurus-RM-7b 82.8 65.9 68.3 72.3
Internlm2-20b-reward 90.2 68.3 62.9 73.6
Skywork-Reward-Gemma-2-27B 93.8 67.3 60.2 73.8
Internlm2-7b-reward 87.6 67.1 67.1 73.9
ArmoRM-Llama3-8B-v0.1 90.4 67.7 64.6 74.2
Nemotron-4-340B-Reward 92.0 69.5 69.9 77.1
Skywork-Reward-Llama-3.1-8B 92.5 70.1 69.3 77.5
INF-ORM-Llama3.1-70B 95.1 70.9 70.5 78.8
GenRMs
Claude-3-5-sonnet-20240620 84.2 61.0 70.6 71.9
Llama3.1-70B-Instruct 84.0 65.5 68.9 72.8
Gemini-1.5-pro 88.2 75.2 56.5 73.3
Skywork-Critic-Llama-3.1-70B 93.3 71.9 65.5 76.9
GPT-4o-0806 86.7 72.5 73.8 77.7
ReasRMs
JudgeLRM 75.2 64.7 53.1 64.3
DeepSeek-PairRM-27B 87.1 58.2
DeepSeek-GRM-27B-RFT 84.5 67.0
DeepSeek-GRM-27B 86.0 69.0
Self-taught-evaluator-llama3.1-70B 90.2 71.4 67.0 76.2
Our Methods
RM-R1-DeepSeek-Distilled-Qwen-7B 80.1 72.4 55.1 69.2
RM-R1-Qwen-Instruct-7B 85.2 70.2 66.4 73.9
RM-R1-Qwen-Instruct-14B 88.2 76.1 69.2 77.8
RM-R1-DeepSeek-Distilled-Qwen-14B 88.9 81.5 68.5 79.6
RM-R1-Qwen-Instruct-32B 91.4 79.1 73.0 81.2
RM-R1-DeepSeek-Distilled-Qwen-32B 90.9 83.9 69.8 81.5
3.2 Main Results
Table 1 compares the overall performance of RM-R1 with existing strongest baseline models. The more detailed numbers on RewardBench, RM-Bench, and RMB are in Table 6,  Table 7, and Table 8 in Appendix F. For the baselines, we reproduce the numbers if essential resources are open-sourced (e.g., model checkpoints, system prompts). Otherwise, we use the numbers reported in the corresponding tech report or benchmark leaderboard. For each benchmark, we select the best-performing models in each category for brevity. Our key findings are summarized below:
State-of-the-Art Performance. On average, our RM-R1-DeepSeek-Distilled-Qwen-14B model surpasses all previous leading Reward Models (RMs), including INF-ORM-Llama3.1-70B, Nemotron-4-340B-Reward, and GPT-4o, while operating at a much smaller scale. Our 32B models, RM-R1-Qwen-Instruct-32B and RM-R1-DeepSeek-Distilled-Qwen-32B, further extend this lead by a notable margin. The success of RM-R1 is attributable to both our meticulously designed training methodology and the effective scaling of our models, as extensively analyzed in Section 4.1 and Section 4.2. In particular, RM-R1 outperforms existing top-tier ScalarRMs. This highlights the considerable potential of ReasRMs, a category where prior GenRMs have exhibited suboptimal performance and are generally not comparable to their scalar counterparts. In contrast to our structured rollout and distillation with RLVR training strategy, prior critique-based methods have relied heavily on rejection sampling and unstructured, self-generated chain-of-thought (CoT) reasoning from instruct models [22, 35], limiting their reasoning capabilities and leading to inferior performance compared to ScalarRMs. Simultaneously, our comprehensive evaluation indicates that the top-performing scalar models on RewardBench do not consistently achieve state-of-the-art (SOTA) performance; in fact, larger models frequently underperform smaller ones. This evaluation underscores the need for a more comprehensive and systematic approach to RM assessment.
Effective Training towards Reasoning for Reward Modeling. Our specialized, reasoning-oriented training pipeline delivers substantial performance gains. For instance, RM-R1-Qwen-Instruct-14B consistently surpasses Self-taught-evaluator-llama-3.1-70B, a reasoning model five times its size. The RM-R1 model series also demonstrates impressive results on RM-Bench, exceeding the top-performing baseline by up to 8.7%. On this most reasoning-intensive benchmark, RM-R1-DeepSeek-Distilled-Qwen-32B establishes a new state-of-the-art. It achieves 91.8% accuracy in math and 74.1% in code, outperforming the previous best models (73% in math and 63% in code) by significant margins. Furthermore, it also records the strongest reasoning performance among our released models on RewardBench. Despite its performance, our Instruct-based models are remarkably data-efficient, reaching competitive performance using only 8.7K examples for distillation—compared to the 800K examples used in training DeepSeek-Distilled [12]. Overall, our study underscores the significant potential of directly adapting large reasoning models into highly effective reward models.
4 Analysis
In this section, we present a series of empirical analyses to understand the key ingredients for training effective reasoning reward models. Our analysis spans scaling effects, design decisions, reasoning ablations, and a case study. We also present additional analysis on training dynamics in Section G.2.
4.1 Training Recipes
We first investigate the key ingredients underlying the successful training of RM-R1. Through a series of ablation studies, we examine our design choices to identify effective strategies for training high-quality reasoning reward models. We compare the following settings: Cold Start RL, Cold Start RL + Rubrics, Cold Start RL + Rubrics + Query Categorization (QC), and Distilled + RL + Rubrics + QC (i.e., RM-R1). The details of these settings are in Section G.1.
Table 2: Ablation study of the design choices for Reasoning Training on RewardBench.
Method Chat Chat Hard Safety Reasoning Average
Instruct (Original) 95.8 74.3 86.8 86.3 85.8
Instruct + Cold Start RL 92.5 81.5 89.7 94.4 89.5
Instruct + Cold Start RL + Rubrics 93.0 82.5 90.8 94.2 90.1
Instruct + Cold Start RL + Rubrics + QC 92.3 82.6 91.6 96.3 90.8
RM-R1 95.3 83.1 91.9 95.2 91.4
In Table 2, we present the results of the ablation studies described above, using the Qwen-2.5-Instruct-32B model as the Instruct (Original) model. Several key conclusions emerge:
  • RL training alone is insufficient. While Cold Start RL slightly improves performance on hard chat and reasoning tasks, it fails to close the gap to fully optimized models.
  • CoR prompting optimizes RM rollout and boosts reasoning performance. Instructing RM-R1 to self-generate chat rubrics or problem solutions before judgment helps overall performance, especially for chat and safety tasks. Incorporating explicit query categorization into the prompt notably improves reasoning performance, suggesting that clearer task guidance benefits learning.
  • Distillation further enhances performance across all axes. Seeding the model with high-quality reasoning traces before RL yields the strongest results, with improvements observed on both hard tasks and safety-sensitive tasks.
[…]
4.2 Scaling Effects
We then investigate how model performance varies with scale, considering both model size and inference-time compute. In some cases – such as ScalarRMs from InternLM2 [6] and Skywork [20] – the smaller models (7B/8B) outperforms the larger ones (20B/27B), showing no advantage of scaling. In this subsection, we show that this trend does not hold for RM-R1, where scaling brings clear and substantial improvements.
4.2.1 Model Sizes
We first analyze the impact of model scale. Our study is based on the Qwen-2.5-Instruct model family at three sizes: 7B, 14B, and 32B. We evaluate performance improvements resulting from our training procedure described in Section 2, with results averaged across three key benchmarks: RewardBench, RM-Bench, and RMB.
For each model size, we compare the original and post-training performance. Figure 4(a) plots the relative improvement (%) with respect to model size. Observing an approximately linear trend, we fit a linear regression model and extrapolate to hypothetical scales of 3B and 72B, shown using faint markers and dashed extensions. The results strongly support a scaling law for reasoning reward models: larger models not only result in an absolute better final performance but also consistently yield greater performance gains. This aligns with the intuition that our training effectively leverages the superior reasoning capabilities of larger models.
(a) Model Size
(b) Inference Compute
Figure 4: Scaling effect of RM-R1. (a) Larger models benefit more from reasoning training. (b) Longer reasoning chains improve RM performance.
Method RewardBench RM-Bench RMB Avg.
Train on Full Data
Instruct + SFT 90.9 75.4 65.9 77.4
Instruct + Distilled + SFT 91.2 76.7 65.4 77.8
RM-R1 * 91.4 79.1 73.0 81.2
Train on 9k (Distillation) Data
Instruct + SFT 88.8 74.8 66.9 76.6
Instruct + Distilled * 89.0 76.3 72.0 79.2
Table 3: Comparison of reasoning-based training versus SFT across benchmarks. * indicates reasoning-based methods. Reasoning training consistently yields better performance.
4.2.2 Inference-time Computation
Next, we examine how model performance varies with different compute budgets measured in number of tokens allowed during inference. Since this is particularly relevant to reasoning-focused models, we fix our base model to DeepSeek-R1-Distill-Qwen-14B. We evaluate average performance across the three key benchmarks using a wide range of inference-time compute budgets: 512, 1024, 2048, 4096, and 8192 tokens.
To ensure a fair comparison, we match the training rollout budget to the inference budget in each setting (i.e., we allow a maximum of $k$ tokens during training for a compute budget of $k$ at inference). All models are trained using GRPO with identical datasets and hyperparameter configurations. Figure 4(b) shows the relationship between compute budget and performance. We observe a clear improvement trend as the inference budget increases. This highlights the benefits of long reasoning chains in reward modeling.
[…]
4.3 Effectiveness of Reasoning Training
We now analyze the impact of reasoning-based training. Here, we demonstrate that reasoning-based training can outperform answer-only approaches. We consider the following settings:
Instruct + SFT. This approach fine-tunes the instruct model directly toward producing the correct final answer using the full dataset, without providing any intermediate reasoning chains.
Instruct + Distilled + SFT. This approach applies SFT (with respect to the answer directly) after the distillation stage, serving as a direct comparison point with RM-R1 trained with RL.
Instruct + RM-R1 (Distilled + RL). This is the full approach proposed in this paper, following the procedure detailed in Section 2.
Instruct + Distilled. This setting uses the model checkpoint immediately after the distillation stage, before any RL fine-tuning.
In summary, methods with “+ RM-R1” or “+ Distilled” represent reasoning-based approaches, while the remaining methods are purely non-reasoning-based approaches. In Table 3, we report the results measured across the three benchmarks. The findings clearly demonstrate that reasoning training significantly benefits reward model performance. Under fair comparisons (i.e., training on exactly the same amount of data), reasoning-based models consistently outperform their SFT-only counterparts. In particular, even high-quality distillation alone, applied to a small subset of the data, provides notable gains, highlighting the value of incorporating structured intermediate reasoning.
[…]
4.4 Case Study
Table 4: Case Study of RM-R1(succinct version). Complete Chatbots and model responses are shown in Table 910, and 11. Cold-Start RL evaluation tends to focus on superficial features (e.g., number of suggestions, relevance), whereas RM-R1 produces high-quality rubrics (e.g., accuracy for factual questions), faithfully follows them, and evaluates responses based on their actual content.
[…]
To gain deeper insights into RM-R1, we further conduct a case study using Qwen-14B-Instruct, comparing its behavior with a cold-start RL model. The results, summarized in Table 4, reveal the following key observations:
Transparent Judging Process.
The reasoning traces generated by RM-R1 are highly interpretable and coherent, reflecting the model’s perception of human preferences. It explicitly articulates why certain responses are better, providing transparency into its evaluation process.
High-Quality, Question-Dependent Rubrics.
RM-R1 accurately understands the question and the context of comparison, correctly prioritizing “accuracy” as the most critical rubric for medical-related questions. In contrast, the cold-start RL model often overlooks the most important factors and instead emphasizes superficial or broadly defined features (e.g., relevance) that are less discriminative. The ability to generate high-quality, question-specific rubrics stems from the knowledge acquired during the distillation stage.
Faithful Adherence to Rubrics and Content-Based Judgement.
RM-R1 grounds its evaluation in the actual content of the chatbot responses. For example, it correctly identifies inaccuracies in Chatbot A’s response based on factual content rather than surface presentation. Furthermore, it systematically evaluates different aspects of the rubric, leading to a structured, interpretable, and verifiable judging process.
5 Related Work
Reward Models (RMs).
Early RMs were typically outcome-focused: trained to predict human preference rankings for complete outputs [42]. Recent advances have looked at providing process supervision, which rewards or evaluates the steps of a model’s reasoning rather than only the final answer. A series of works propose to train process reward models that judge the correctness of intermediate reasoning steps [19, 10, 27]. A limitation of many PRMs is their heavy reliance on curated step-level human labels or specific schemas, and they often remain domain-specific. Zhang et al. [40] propose Generative Verifiers, framing reward modeling as a next-token prediction task. This allows the reward model to leverage chain-of-thought and even use majority voting over multiple sampled rationales to make more reliable judgments. DeepSeek-GRM [22] and JudgeLRM [7] have studied using reasoning models as generative reward models, which are the most relevant research to ours. However, their main focus is on the effect of scaling inference-time computation on reward modeling. On the contrary, our work is the first to provide a systematic empirical comparison of different reward model training paradigms, shedding light on when and why a distilled and RL-trained reward model like RM-R1 has advantages over the conventional approaches.
Reinforcement Learning from Human Feedback (RLHF).
Early works [8] first demonstrated that reinforcement learning could optimize policies using a reward model trained from human pairwise preferences. Subsequent studies applied RLHF to large-scale language models using policy optimization algorithms such as PPO [26]. For example,  Ziegler et al. [45] fine-tuned GPT-2 via PPO on human preference rewards, and Stiennon et al. [32] showed that RLHF could significantly improve the quality of summarization by optimizing against a learned preference model. More recently,  Ouyang et al. [24] used a similar PPO-based pipeline to train InstructGPT, establishing the modern RLHF paradigm for instruction-following models. Recently, Verifiable supervision techniques have also emerged: DeepSeek-R1 [12] uses a form of self-verification during RLHF to reward correct reasoning steps, rather than only final-answer quality. This method incentivizes policies to produce outputs that can be verified for correctness, bridging the gap between pure preference-based feedback and ground-truth signals. However, even with such innovations, most RLHF implementations still treat reward modeling and reasoning as separate stages.
6 Conclusion and Future Work
In this paper, we revisited reward modeling through the lens of reasoning. We introduced RM-R1, a family of ReasRMs that effectively generate explicit chains of rubrics and rationales, and scale with both model size and inference compute. Across three public benchmarks, RM-R1 matched or surpassed commercial and open-source RMs while producing more interpretable judgments. Ablation investigations reveal that (1) task-type categorization, (2) bootstrapping from high-quality reasoning traces, and (3) RL fine-tuning are all indispensable. Qualitative analyses further showed that RM-R1 learns to prioritize high-impact rubrics, faithfully follow its own criteria and justify coherently. Future work includes active preference collection, where ReasRMs use active learning to query human preference only when the current rubric set is insufficient for a new preference sample. Finally, it would be natural to extend our study to multimodal/agentic reward modeling scenarios.
Acknowledgments and Disclosure of Funding
This research is based upon work supported DARPA ITM Program No. FA8650-23-C-7316, and the AI Research Institutes program by National Science Foundation and the Institute of Education Sciences, U.S. Department of Education through Award # 2229873 - AI Institute for Transforming Education for Children with Speech and Language Processing Challenges. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
Appendix A User Prompt for DeepSeek-Distilled Reasoning Models
Large reasoning models such as DeepSeek-R1-distilled models [12] do not have a system prompt, so we show the user prompt for rollouts in Figure 5.
Chain-of-Rubrics (CoR) Rollout for Reasoning Models Please act as an impartial judge and evaluate the quality of the responses provided by two AI Chatbots to the Client question displayed below. … [Pairwise Input Content] … Output your final verdict at last by strictly following this format: ’<answer>[[A]]</answer>’ if Chatbot A is better, or ’<answer>[[B]]</answer>’ if Chatbot B is better.
Figure 5: The user prompt used for RM-R1 rollout (for reasoning models).
Appendix B Details of Reasoning Chain Generation
We now expand on the details of generating high-quality reasoning chains. We first use the same prompt to query Claude-3.7-Sonnet, generating initial reasoning traces. However, approximately 25% of these traces are incorrect, primarily on harder chat tasks. To correct these cases, we pass the original prompt, the incorrect trace, and the correct final answer to OpenAI-O3, which then generates a corrected reasoning trace aligned with the right answer.
This two-stage process yields a high-quality distillation set. We deliberately choose the order—first Claude, then O3—based on qualitative observations: Claude excels at solving easier tasks and maintaining attention to safety considerations, whereas O3 performs better on harder tasks but tends to overemphasize helpfulness at the expense of safety. We select approximately 12% of the training data (slightly fewer than 9K examples) for distillation. This is then followed by RL training.
Appendix C Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO) [28] is a variant of Proximal Policy Optimization (PPO) [26], which obviates the need for additional value function approximation, and uses the average reward of multiple sampled outputs produced in response to the same prompt as the baseline. More specifically, for each prompt $x$, GRPO samples a group of outputs $\{y_{1},y_{2},\cdots,y_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy model by maximizing the following objective:
$\begin{split}\mathcal{J}_{\text{GRPO}}(\theta)=\,&\mathbb{E}_{x\sim\mathcal{D},\,\{j_{i}\}_{i=1}^{G}\sim r_{\theta_{\text{old}}}(j\mid x)}\bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|j_{i}|}\sum_{t=1}^{|j_{i}|}\Big{\{}\min\Big{(}\frac{r_{\theta}(j_{i,t}\mid x,j_{i,<t})}{r_{\theta_{\text{old}}}(j_{i,t}\mid x,j_{i,<t})}\hat{A}_{i,t},\\[4.0pt] &\text{clip}\big{(}\frac{r_{\theta}(j_{i,t}\mid x,j_{i,<t})}{r_{\theta_{\text{old}}}(j_{i,t}\mid x,j_{i,<t})},1-\epsilon,1+\epsilon\big{)}\hat{A}_{i,t}\Big{)}-\beta\,\mathbb{D}_{\mathrm{KL}}\left[r_{\theta}(\cdot\mid x)\,\|\,\pi_{\text{ref}}(\cdot\mid x)\right]\Big{\}}\bigg{]},\end{split}$ (9)
where $\beta$ is a hyperparameter balancing the task specific loss and the KL-divergence. Specifically, $\hat{A}_{i}$ is computed using the rewards of a group of responses within each group $\{r_{1},r_{2},\ldots,r_{G}\}$, and is given by the following equation:
$\hat{A}_{i}=\frac{r_{i}-{\text{mean}(\{r_{1},r_{2},\cdots,r_{G}\})}}{{\text{std}(\{r_{1},r_{2},\cdots,r_{G}\})}}.$ (10)
Appendix D Experiment Setups
D.1 Benchmarks
In this paper, we consider the following three benchmarks:
RewardBench [17]: RewardBench is one of the first endeavors towards benchmarking reward models through prompt-chosen-rejected trios, covering four categories: chat, chat-hard, reasoning, and safety, with 358, 456, 740, and 1431 samples, respectively.
RM-Bench [21]: Building on RewardBench, RM-Bench evaluates reward models for their sensitivity to subtle content differences and robustness against style biases. It includes four categories: Chat, Safety, Math, and Code, with 129, 441, 529, and 228 samples, respectively. Each sample contains three prompts of varying difficulty. RM-Bench is the most reasoning-intensive benchmark among those we consider.
RMB [43]: Compared with RewardBench and RM-Bench, RMB offers a more comprehensive evaluation of helpfulness and harmlessness. It includes over 49 real-world scenarios and supports both pairwise and Best-of-N (BoN) evaluation formats. RMB comprises 25,845 instances in total—37 scenarios under the helpfulness alignment objective and 12 under harmlessness.
D.2 Preference Datasets
We consider the following datasets for training:
Skywork Reward Preference 80K [20] is a high-quality collection of pairwise preference data drawn from a variety of domains, including chat, safety, mathematics, and code. It employs an advanced data filtering technique to ensure preference reliability across tasks. However, we identify a notable issue with this dataset: all samples from the magpie_ultra source exhibit a strong spurious correlation, where rejected responses consistently contain the token “<im_start>,” while accepted responses do not. Additionally, responses from this source show a systematic bias—accepted responses are typically single-turn, while rejected responses are multi-turn. This problematic subset constitutes approximately 30% of the Skywork dataset and primarily covers mathematics and code domains. To avoid introducing spurious correlations into training, we exclude all magpie_ultra data and retain only the cleaned subset for our experiments.
Code-Preference-Pairs is a high-quality coding preference dataset. It is constructed by prompting a model with original code, introducing deliberate bugs, and manipulating examples (e.g., swapping broken and corrected versions, removing error comments) to generate fine-grained preference pairs. We subsample 8K examples from this dataset for use in our experiments.
Math-DPO-10K [16] is a high-quality stepwise preference dataset focused on mathematical reasoning. We use the full dataset in our experiments.
A global statistics of our training dataset is summarized in Table 5.
Table 5: Global Statistics of our Training Dataset. * indicates the source is from Skywork-Reward-Preference-80K-v0.2.
Source Size Domain
magpile_pro_llama3.1* 29682 Reasoning
offset_bias* 8504 Chat (length bias)
helpsteer2* 7221 Chat
wildguard* 6709 Safety
magpile_pro* 2030 Chat
Code-Preference-Pairs 8000 Reasoning
Math-DPO-10K 10000 Reasoning
D.3 Baselines
We compare RM-R1 with RMs from three categories:
ScalarRMs. ScalarRMs produce a score for model response directly, predicting preference through single numeric scores without explicit reasoning traces. This category includes models such as Eurus-RM [39], Internlm2 [6] SteerLM-RM [36], Nemotron-RM [2], Tulu-v2.5 [15], Starling-RM [44], ArmoRM [34], Skywork-RM [20], etc. While these models often achieve strong results on well-defined benchmarks, they generally lack interpretability and struggle to capture fine-grained reasoning.
GenRMs. Generative reward models (GenRMs) offer more expressive feedback by producing free-form textual judgments, typically without further training. This includes the widely used LLM-as-a-Judge setup [41], where pretrained language models are prompted to explain and evaluate responses. We also categorize under GenRMs models that directly generate output answers without intermediate reasoning steps. Representative examples include LLaMA [11], Qwen [37], Claude [3], GPT-4o [1, 14], Gemini 1.5 Pro [25], and Skywork-Critic [30]. By leveraging LLMs’ generative capabilities, these models enhance interpretability through natural language rationales and explanations.
ReasRMs. Reasoning-enhanced reward models (ReasRMs) explicitly incorporate reasoning processes before their final judgments, often trained through critiques or chain-of-thought strategies. Notable examples are JudgeLRM [7], Critique-RM [38], DeepSeek-GRM [22], Self-taught Evaluators [35] and our proposed RM-R1 models. These models excel in tasks demanding rigorous reasoning, safety evaluations, and nuanced preference judgments due to their grounding in structured critical thinking.
Appendix E Implementation Details
Our training framework is based on VERL [29] and OpenRLHF [13]. For Instruct models, we use 8.7k data for distillation and 64k for RLVR. For Deepseek-Distilled models, we use the full data for RLVR.
[…]
RLVR Stage. We use the VERL framework for all GRPO training. The training batch size is fixed at 1024, with a mini-batch size of 128. We adopt Fully Sharded Data Parallel (FSDP) to improve memory efficiency. For rollout generation, we use vLLM with tensor parallelism size 4 and GPU memory utilization capped at 0.4. Sampling follows default parameters (temperature = 1.0, top-p = 1.0). KL regularization is applied with a coefficient of $1\mathrm{e}{-3}$ and a clip ratio of 0.2. Each prompt is sampled with 7 candidate responses.
The maximum input sequence length is 4,096 tokens, and the maximum response length is 8,192 tokens. Learning rates are set separately for the two model variants:
[…]
Appendix F Full Experiment Result
In this section, we provide the full experiment results and a more comprehensive coverage of existing baselines. The results of RewardBench, RM-Bench, and RMB are provided in Table 6Table 7Table 8, respectively.
[…]
Models Chat Chat_Hard Safety Reasoning Overall
ScalarRMs
Eurus-RM-7b 98.0 65.6 81.4 86.3 82.8
Internlm2-7b-reward 99.2 69.5 87.2 94.5 87.6
SteerLM-RM 70B 91.3 80.3 92.8 90.6 88.8
Cohere-0514 96.4 71.3 92.3 97.7 89.4
Internlm2-20b-reward 98.9 76.5 89.5 95.8 90.2
ArmoRM-Llama3-8B-v0.1 96.9 76.8 90.5 97.3 90.4
Nemotron-4-340B-Reward 95.8 87.1 91.5 93.6 92.0
[…]
INF-ORM-Llama3.1-70B 96.6 91.0 93.6 99.1 95.1
GenRMs
Llama3.1-8B-Instruct 85.5 48.5 75.6 72.1 70.4
Prometheus-8*7B-v2 93.0 47.1 80.5 77.4 74.5
Llama3.1-70B-Instruct 97.2 70.2 82.8 86.0 84.0
Llama3.1-405B-Instruct 97.2 74.6 77.6 87.1 84.1
Claude-3-5-sonnet-20240620 96.4 74.0 81.6 84.7 84.2
GPT-4o-0806 96.1 76.1 86.6 88.1 86.7
Gemini-1.5-pro 92.3 80.6 87.9 92.0 88.2
SFR-LLaMa-3.1-70B-Judge-r 96.9 84.8 91.6 97.6 92.7
[…]
ReasRMs
JudgeLRM 92.9 56.4 78.2 73.6 75.2
SynRM 38.0 82.5 74.1 87.1 70.4
RM-R1-DeepSeek-Distilled-Qwen-7B 88.9 66.2 78.4 87.0 80.1
CLoud 97.0 58.0 84.0 92.0 82.8
DeepSeek-GRM-16B 90.8 74.3 84.7 81.8 82.9
DeepSeek-GRM-27B-RFT 94.7 77.2 87.0 79.2 84.5
RM-R1-Qwen-Instruct-7B 94.1 74.6 85.2 86.7 85.2
DeepSeek-GRM-27B 94.1 78.3 88.0 83.8 86.0
DeepSeek-PairRM-27B 95.5 86.8 52.3 92.0 87.1
RM-R1-Qwen-Instruct-14B 93.6 80.5 86.9 92.0 88.2
RM-R1-DeepSeek-Distilled-Qwen-14B 91.3 79.4 89.3 95.5 88.9
Self-taught-evaluator-llama3.1-70B 96.9 85.1 89.6 88.4 90.0
RM-R1-DeepSeek-Distilled-Qwen-32B 95.3 80.3 91.1 96.8 90.9
RM-R1-Qwen-Instruct-32B 95.3 83.1 91.9 95.2 91.4
Table 7: The full results of tested reward models on RM-Bench. Chat, Math, Code, Safety show the model’s Average Accuracy on each domain. Easy, Normal, Hard show the model’s Accuracy on each difficulty level across all domains. Bold numbers indicate the best performance, Underlined numbers indicate the second best.
Models Chat Math Code Safety Easy Normal Hard Avg
ScalarRMs
steerlm-70b 56.4 53.0 49.3 51.2 48.3 54.9 54.3 52.5
tulu-v2.5-70b-preference-mix-rm 58.2 51.4 55.5 87.1 72.8 65.6 50.7 63.0
Mistral-7B-instruct-Unified-Feedback 56.5 58.0 51.7 86.8 87.1 67.3 35.3 63.2
RM-Mistral-7B 57.4 57.0 52.7 87.2 88.6 67.1 34.9 63.5
Eurus-RM-7b 59.9 60.2 56.9 86.5 87.2 70.2 40.2 65.9
internlm2-7b-reward 61.7 71.4 49.7 85.5 85.4 70.7 45.1 67.1
Skywork-Reward-Gemma-2-27B 69.5 54.7 53.2 91.9 78.0 69.2 54.9 67.3
ArmoRM-Llama3-8B-v0.1 67.8 57.5 53.1 92.4 82.2 71.0 49.8 67.7
GRM-llama3-8B-sftreg 62.7 62.5 57.8 90.0 83.5 72.7 48.6 68.2
internlm2-20b-reward 63.1 66.8 56.7 86.5 82.6 71.6 50.7 68.3
Llama-3-OffsetBias-RM-8B 71.3 61.9 53.2 89.6 84.6 72.2 50.2 69.0
Nemotron-340B-Reward 71.2 59.8 59.4 87.5 81.0 71.4 56.1 69.5
URM-LLaMa-3.1-8B 71.2 61.8 54.1 93.1 84.0 73.2 53.0 70.0
Skywork-Reward-Llama-3.1-8B 69.5 60.6 54.5 95.7 89.0 74.7 46.6 70.1
INF-ORM-Llama3.1-70B 66.3 65.6 56.8 94.8 91.8 76.1 44.8 70.9
GenRMs
tulu-v2.5-dpo-13b-chatbot-arena-2023 64.9 52.3 50.5 62.3 82.8 60.2 29.5 57.5
tulu-v2.5-dpo-13b-nectar-60k 56.3 52.4 52.6 73.8 86.7 64.3 25.4 58.8
stablelm-2-12b-chat 67.2 54.9 51.6 65.2 69.1 63.5 46.6 59.7
tulu-v2.5-dpo-13b-stackexchange-60k 66.4 49.9 54.2 69.0 79.5 63.0 37.2 59.9
Nous-Hermes-2-Mistral-7B-DPO 58.8 55.6 51.3 73.9 69.5 61.1 49.1 59.9
Claude-3-5-sonnet-20240620 62.5 62.6 54.4 64.4 73.8 63.4 45.9 61.0
tulu-v2.5-dpo-13b-hh-rlhf-60k 68.4 51.1 52.3 76.5 53.6 63.0 69.6 62.1
tulu-2-dpo-13b 66.4 51.4 51.8 85.4 86.9 66.7 37.7 63.8
SOLAR-10.7B-Instruct-v1.0 78.6 52.3 49.6 78.9 57.5 67.6 69.4 64.8
Llama3.1-70B-Instruct 64.3 67.3 47.5 83.0 74.7 67.8 54.1 65.5
Skywork-Critic-Llama-3.1-70B 71.4 64.6 56.8 94.8 85.6 73.7 56.5 71.9
GPT-4o-0806 67.2 67.5 63.6 91.7 83.4 75.6 58.7 72.5
Gemini-1.5-pro 71.6 73.9 63.7 91.3 83.1 77.6 64.7 75.2
ReasRMs
JudgeLRM 59.9 59.9 51.9 87.3 73.2 766.2 54.8 64.7
RM-R1-Qwen-Instruct-7B 66.6 67.0 54.6 92.6 79.2 71.7 59.7 70.2
Self-taught-evaluator-llama3.1-70B 73.4 65.7 56.3 90.4 80.2 74.5 59.7 71.5
RM-R1-DeepSeek-Distilled-Qwen-7B 64.0 83.9 56.2 85.3 75.9 73.1 68.1 72.4
RM-R1-Qwen-Instruct-14B 75.6 75.4 60.6 93.6 82.6 77.5 68.8 76.1
RM-R1-Qwen-Instruct-32B 75.3 80.2 66.8 93.9 86.3 80.5 70.4 79.1
RM-R1-DeepSeek-Distilled-Qwen-14B 71.8 90.5 69.5 94.1 86.2 83.6 74.4 81.5
RM-R1-DeepSeek-Distilled-Qwen-32B 74.2 91.8 74.1 95.4 89.5 85.4 76.7 83.9
Table 8: The leaderboard of RMB, ranked by the average score of all subsets. Bold numbers indicate the best performance, Underlined numbers indicate the second best.
Helpfulness Harmlessness
Models BoN Pairwise BoN Pairwise Overall
ScalarRMs
Tulu-v2.5-13b-preference-mix-rm 0.355 0.562 0.351 0.545 0.453
SteerLM-RM 70B 0.502 0.574 0.578 0.673 0.582
Skywork-Reward-Gemma-2-27B 0.472 0.653 0.561 0.721 0.602
Internlm2-20b-reward 0.585 0.763 0.499 0.670 0.629
ArmoRM-Llama3-8B-v0.1 0.636 0.787 0.497 0.663 0.646
Internlm2-7b-reward 0.626 0.782 0.563 0.712 0.671
Eurus-RM-7b 0.679 0.818 0.543 0.693 0.683
Skywork-Reward-Llama-3.1-8B 0.627 0.781 0.603 0.759 0.693
INF-ORM-Llama3.1-70B 0.650 0.798 0.607 0.767 0.705
Starling-RM-34B 0.604 0.774 0.674 0.795 0.712
GenRMs
Llama2-70b-chat 0.289 0.613 0.249 0.602 0.438
Llama3.1-8B-Instruct 0.365 0.675 0.267 0.653 0.490
Gemini-1.5-pro 0.536 0.763 0.299 0.661 0.565
Mixtral-8x7B-Instruct-v0.1 0.480 0.706 0.491 0.671 0.587
skywork-critic-llama3.1-8B 0.600 0.725 0.578 0.578 0.620
skywork-critic-llama3.1-70B 0.640 0.753 0.614 0.614 0.655
Llama3.1-70B-Instruct 0.648 0.811 0.558 0.739 0.689
Mistral-Large-2407 0.678 0.817 0.583 0.725 0.701
Claude-3-5-sonnet 0.705 0.838 0.518 0.764 0.706
Qwen2-72B-Instruct 0.645 0.810 0.649 0.789 0.723
GPT-4o-2024-05-13 0.639 0.815 0.682 0.814 0.738
ReasRMs
JudgeLRM 0.363 0.699 0.363 0.674 0.531
RM-R1-DeepSeek-Distilled-Qwen-7B 0.451 0.658 0.429 0.664 0.551
RM-R1-Qwen-Instruct-7B 0.543 0.740 0.608 0.765 0.664
Self-taught-evaluator-llama3.1-70B 0.616 0.786 0.546 0.733 0.670
Deepseek-GRM-27B-RFT 0.592 0.801 0.548 0.765 0.670
RM-R1-DeepSeek-Distilled-Qwen-14B 0.593 0.765 0.613 0.769 0.685
Deepseek-GRM-27B 0.623 0.805 0.570 0.761 0.690
RM-R1-Qwen-Instruct-14B 0.594 0.776 0.620 0.778 0.692
RM-R1-DeepSeek-Distilled-Qwen-32B 0.620 0.782 0.618 0.771 0.698
RM-R1-Qwen-Instruct-32B 0.636 0.791 0.682 0.809 0.730
Table 9: Complete questions and chatbot responses.
Question: symptoms of sickle-cell disease Chatbot A: 1. Painful red or yellow skin lesions 2. Frequent infections 3. Fatigue and weakness 4. Anemia 5. Slower growth and development 6. Jaundice 7. Swelling of hands and feet 8. Vision loss 9. Tiredness and weakness after physical activity 10. Breathlessness 11. Fever, chills, and night sweats. Chatbot B Sickle cell disease is a genetic disorder that affects the shape of red blood cells. It is most common in African Americans, but can also affect people from other ethnic groups. The symptoms of sickle cell disease can vary widely from person to person, but some common symptoms include: 1. Fatigue: People with sickle cell disease often experience fatigue and weakness, even when they have not been physically active. 2. Pain: Sickle cell disease can cause pain in the body, particularly in the arms, legs, back, and abdomen. Pain can be mild or severe and can last for days or weeks. 3. Fever: People with sickle cell disease may be prone to infections and may experience frequent fevers. 4. Anemia: Sickle cell disease can cause anemia, a condition in which the body does not have enough red blood cells to carry oxygen throughout the body. 5. Delayed growth and development: Children with sickle cell disease may be smaller than their peers due to delayed growth and development. 6. Infections: People with sickle cell disease are at increased risk for infections, particularly in the bloodstream and urinary tract. 7. Pigmentation: Sickle cell disease can cause dark areas of pigmentation in the body, particularly on the hands and feet. It’s important to note that not everyone with sickle cell disease will experience all of these symptoms, and some people may not experience any symptoms at all. Therefore, it’s important to consult with a healthcare provider if you suspect you or someone you know may have sickle cell disease.
Table 10: Cold-Start RL Judges. The rubrics and evaluation mainly center around superficial shallow features (e.g., number of points listed) without actually taking their accuracy into account.
Cold-start RL evaluation <type>Chat</type> <rubric> The evaluation criteria for the responses are as follows: - <item>Relevance to the question</item> (40%): The response should directly address the symptoms of sickle cell disease. - <item>Comprehensiveness</item> (30%): The response should include a wide range of symptoms commonly associated with sickle cell disease. - <item>Clarity and organization</item> (30%): The response should be easy to read and understand, with symptoms listed in a clear and organized manner. <justify> The question specifically asks for the symptoms of sickle cell disease, so the responses should be directly relevant to this topic. Comprehensiveness is important because sickle cell disease can have a variety of symptoms, and a comprehensive list will provide a more complete picture of the condition. Clarity and organization are important because the symptoms can be complex, and a well-organized response will make it easier for the client to understand the information. </justify> </rubric> <eval> <quote_A> 1. Painful red or yellow skin lesions 2. Frequent infections 3. Fatigue and weakness 4. Anemia 5. Slower growth and development 6. Jaundice 7. Swelling of hands and feet 8. Vision loss 9. Tiredness and weakness after physical activity 10. Breathlessness 11. Fever, chills, and night sweats. </quote_A> <summary_A> Chatbot A provided a list of symptoms that are directly relevant to sickle cell disease. The response is comprehensive and includes a wide range of symptoms commonly associated with the condition. The symptoms are listed in a clear and organized manner, making it easy for the client to understand the information. </summary_A> <quote_B> Sickle cell disease is a genetic disorder that affects the shape of red blood cells. It is most common in African Americans, but can also affect people from other ethnic groups. The symptoms of sickle cell disease can vary widely from person to person, but some common symptoms include: 1. Fatigue: People with sickle cell disease often experience fatigue and weakness, even when they have not been physically active. 2. Pain: Sickle cell disease can cause pain in the body, particularly in the arms, legs, back, and abdomen. Pain can be mild or severe and can last for days or weeks. 3. Fever: People with sickle cell disease may be prone to infections and may experience frequent fevers. 4. Anemia: Sickle cell disease can cause anemia, a condition in which the body does not have enough red blood cells to carry oxygen throughout the body. 5. Delayed growth and development: Children with sickle cell disease may be smaller than their peers due to delayed growth and development. 6. Infections: People with sickle cell disease are at increased risk for infections, particularly in the bloodstream and urinary tract. 7. Pigmentation: Sickle cell disease can cause dark areas of pigmentation in the body, particularly on the hands and feet. </quote_B> <summary_B> Chatbot B provided a detailed explanation of sickle cell disease and a list of symptoms. The response is relevant to the question and includes a wide range of symptoms commonly associated with the condition. However, the response is less organized than Chatbot A’s response, and the explanation of sickle cell disease may be unnecessary for the client’s question. </summary_B> </eval> <answer>[[A]]</answer>
Table 11: RM-R1 RL Judges. The rubrics and evaluation are well-designed and reasonable, assessing the effectiveness of the model response.
Warm-start RL (RM-R1) evaluation <type>Chat</type> <rubric> 1. Accuracy of information (40 %): How correctly does the response describe the symptoms of sickle-cell disease? 2. Comprehensiveness (30 %): How thoroughly does the response cover the range of symptoms associated with sickle-cell disease? 3. Clarity and organization (20 %): How well-structured and easy to understand is the information presented? 4. Helpfulness for the user (10 %): How useful is the information for someone seeking to understand sickle-cell disease symptoms? <justify> These criteria focus on evaluating how well each response addresses the user’s question about sickle-cell disease symptoms. Accuracy is weighted highest because providing correct information about medical conditions is crucial. Comprehensiveness is important, as sickle-cell disease has multiple symptoms that vary in severity and presentation. Clarity and organization affect how easily the user can understand the information. Helpfulness considers how practical and actionable the information is for the user. </justify> </rubric> <eval> Comparing the two responses: Accuracy of information: <quote_A>1. Painful red or yellow skin lesions 2. Frequent infections 3. Fatigue and weakness 4. Anemia 5. Slower growth and development 6. Jaundice 7. Swelling of hands and feet 8. Vision loss 9. Tiredness and weakness after physical activity 10. Breathlessness 11. Fever, chills, and night sweats.</quote_A> <quote_B>1. Fatigue: People with sickle-cell disease often experience fatigue and weakness, even when they have not been physically active. 2. Pain: Sickle-cell disease can cause pain in the body, particularly in the arms, legs, back, and abdomen. Pain can be mild or severe and can last for days or weeks. 3. Fever: People with sickle-cell disease may be prone to infections and may experience frequent fevers. 4. Anemia: Sickle-cell disease can cause anemia, a condition in which the body does not have enough red blood cells to carry oxygen throughout the body. 5. Delayed growth and development: Children with sickle-cell disease may be smaller than their peers due to delayed growth and development. 6. Infections: People with sickle-cell disease are at increased risk for infections, particularly in the bloodstream and urinary tract. 7. Pigmentation: Sickle-cell disease can cause dark areas of pigmentation in the body, particularly on the hands and feet.</quote_B> Chatbot A provides a list of symptoms without explanation, but some listed items (e.g. “painful red or yellow skin lesions,” “vision loss”) are not typical of sickle-cell disease. Chatbot B offers more accurate information, explaining each symptom’s nature and severity. Comprehensiveness: Chatbot A lists 11 symptoms, several of which are inaccurate or uncommon. Chatbot B lists 7 symptoms that are more representative and supplies contextual detail about variability and medical advice. Clarity and organization: Chatbot A’s plain numbered list lacks context, making severity unclear. Chatbot B structures the list with explanations, improving readability and understanding. Helpfulness: Chatbot A is less helpful due to inaccuracies and missing context. Chatbot B is more useful, providing accurate descriptions, context, and guidance. </eval> <answer>[[B]]</answer>
Appendix G Supplementary Information for Section 4
G.1 Ablation Settings
Cold Start RL. This approach generally involves pure RL, with rule-based rewards centered on answer correctness and format compliance. Such strategies have achieved notable success in advanced mathematical problem solving [28].
In this setting, we replicate this conventional training setup. Specifically, we use a combination of a format reward and an answer reward:
$\mathcal{R}_{\text{format}}=\begin{cases}1&\text{ if format matches,}\\ 0&\text{ otherwise,}\end{cases}\quad\text{and}\quad\mathcal{R}_{\text{answer}}=\begin{cases}1&\text{ if answer matches,}\\ 0&\text{ otherwise.}\end{cases}$ (11)
The total reward is the sum $\mathcal{R}=\mathcal{R}_{\text{answer}}+\mathcal{R}_{\text{format}}$. We use the prompt template shown in Figure 7, a version without any guidance on structured reasoning.
Chain -of -Rubrics (CoR) Roll-out for Instruct Models (no categorization of task types) Please act as an impartial judge and evaluate the quality of the responses provided by two AI Chatbots to the Client’s question displayed below.
Instructions
1. Begin your evaluation by generating the rubric criteria tailored to the Client’s question and context.
   Enclose the rubric in <rubric> … </rubric> tags.
2. Assign weights to each rubric item based on their relative importance. 3. Within <rubric>, include a <justify> … </justify> section explaining the rationale behind the chosen criteria and weights. 4. Compare both Chatbot responses using the rubric. 5. Include your evaluation in <eval> … </eval> tags.
   Support your analysis using:
    - <quote_A> … </quote_A> for direct quotes from Chatbot A
    - <summary_A> … </summary_A> for paraphrased summaries of Chatbot A
    - <quote_B> … </quote_B> for direct quotes from Chatbot B
    - <summary_B> … </summary_B> for paraphrased summaries of Chatbot B
6. Conclude with your final judgment using:
    <answer>[[A]]</answer>  or  <answer>[[B]]</answer>
Important Notes:
- Be objective and base your evaluation strictly on the content of the responses.
- Do not let the response order, length, or Chatbot names bias your judgment.
Figure 6: The system prompt of the ablation study on cold start RL without categorization of task types.
Chain -of -Rubrics (CoR) Roll-out for Instruct Models (no rubrics) Please act as an impartial judge and evaluate the quality of the responses provided by two AI Chatbots to the Client’s question displayed below.
You should choose the chatbot that follows the client’s instructions and answers the client’s question better. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the chatbots. Be as objective as possible. First, compare the chatbot responses and provide your evaluations. Then, conclude with your verdict using exactly this format: <answer>[[A]]</answer> if Chatbot A is better, <answer>[[B]</answer> if Chatbot B is better.
Figure 7: The system prompt of the ablation study on cold start RL without any rubrics.
Cold Start RL + Rubrics. To examine the influence of structured reasoning in final model performance, compared with the last setting, we use the prompt template shown in Figure 6. Compared with the last setting, the model is prompted to generate rubrics and evaluate accordingly. However, compared with the final system prompt of RM-R1 Figure 3, all input prompts are treated uniformly—that is, chat and reasoning tasks are not distinguished.
Cold Start RL + Rubrics + Query Categorization (QC). This setting largely follows the previous one, with a key modification: prompting the LM to first categorize the task into reasoning or chat, and then apply different strategies for handling those tasks. Intuitively, reinforcement learning alone can effectively explore reasoning tasks, a domain where it has already achieved considerable success. Here, we incorporate the system prompt shown in Figure 3, which explicitly distinguishes between chat and reasoning tasks.
For reasoning tasks specifically, we note that answer quality is closely tied to correctness, and that high-level rubrics may be less effective than simply evaluating whether the model can solve the problem and verify its own answer. Thus, this setting emphasizes correctness-based evaluation guided by task classification in the prompt.
Distilled + RL + Rubrics + QC (RM-R1). Building on the previous setup, we introduce an additional distillation stage from stronger teacher models as a warm start before RL training. The motivation is that with RL alone, weaker models (especially at smaller scales) often fail to explore high-quality rubrics and convincing reasoning chains for chat tasks throughout the RL training process. Distilling strong reasoning traces on a small subset of data can effectively mitigate this limitation.
(a) Cold Start RL
(b) Warm Start RL
Figure 8: RL training dynamics under different settings: (a) Cold Start RL (Eq. 11) and (b) Warm Start RL (Eq. 8). In Cold Start RL, the response length steadily increases as the model learns to reason, but training becomes unstable near the end. In Warm Start RL, the model exhibits more stable training, with effective refinement of reasoning traces throughout the process.
G.2 Training Dynamics
We analyze the training dynamics of RM-R1 using the Qwen-2.5-14B-Instruct model by tracking both response length and reward progression throughout RL training. We consider two settings: (a) Cold Start RL, and (b) Warm Start RL following reasoning-chain distillation. We present the finding in Figure 8.
In the Cold Start RL setting, we observe that the model gradually learns to reason, as reflected by a steady increase in response length over the course of training. However, training becomes unstable near the end, with a sharp drop in the reward curve, suggesting potential issues such as overfitting.
In contrast, under Warm Start RL, the model begins with stronger initial reasoning abilities, exhibiting longer responses from the outset. Interestingly, the model first learns to produce more concise reasoning traces before gradually increasing response length again as training progresses. The reward curve rises smoothly and consistently throughout training, demonstrating more stable and efficient learning compared to the Cold Start setting.